40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3730     Entry time: Fri Oct 15 21:25:23 2010
Author: Suresh 
Type: Update 
Category: IOO 
Subject: 2W NPRO laser output power drop question 

  The power meter used in the measurements of elog entries 2822, 3698 and 3709 was the Ophir PD300-3W.  This power head has several damaged patches  and a slight movement of the laser spot changes the reading considerably.  To verify I checked the power out with another power meter (the Vector S310) and found that there is no significant variation of the power output with the temperature of the laser.  And the power at 2.1A of diode current is 2W with 10% fluctuation arising from slight repositioning of the laser head.  There are regions of the Ophir PD300 which show the laser power to be about 1.9W.

ELOG V3.1.3-